# Close-Talk, and Pseudo-Label Based Far-Field, Speech Enhancement

### v1

@@ Abstract
The current dominant approach for neural speech enhancement is via purely-supervised deep learning on simulated pairs of far-field noisy-reverberant speech (i.e., mixtures) and clean speech. The trained models, however, often exhibit limited generalizability to real-recorded mixtures. To deal with this, this paper investigates training enhancement models directly on real mixtures. However, a major difficulty challenging this approach is that, since the clean speech of real mixtures is unavailable, there lacks a good supervision for real mixtures. In this context, assuming that a training set consisting of real-recorded pairs of close-talk and far-field mixtures is available, we propose to address this difficulty via close-talk speech enhancement, where an enhancement model is first trained on simulated mixtures to enhance real-recorded close-talk mixtures and the estimated close-talk speech can then be utilized as a supervision (i.e., pseudo-label) for training far-field speech enhancement models directly on the paired real-recorded far-field mixtures. We name the proposed system ctPuLSE. Evaluation results on the CHiME-4 dataset show that ctPuLSE can derive high-quality pseudo-labels and yield far-field speech enhancement models with strong generalizability to real data.
Index Terms:
Close-talk speech enhancement, pseudo-label based speech enhancement, robust automatic speech recognition.
@@ I Introduction
Dramatic progress has been made in speech enhancement [1], thanks to the rapid development of deep learning. The current dominant approach is based on supervised deep learning, where paired noisy-reverberant speech (i.e., mixtures) and clean speech are simulated and utilized to train deep neural networks (DNN) to predict the clean speech based on its paired mixtures in a purely-supervised, discriminative way [1, 2]. The trained models, however, often exhibit limited generalizability to real-recorded mixtures [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13], mainly because the simulated training data is typically, and in many cases inevitably, mismatched with real test data.
@@
Although ctPuLSE is simple and straightforward, it obtains strong robust ASR and speech enhancement performance on the public CHiME-4 dataset [17], the most popular benchmark to date in robust ASR and speech enhancement. The evaluation results suggest that ctPuLSE can effectively train enhancement models on real-recorded far-field mixtures, and can yield better generalizability to real data than purely-supervised models trained only on simulated data. A sound demo is provided in the link below [^1] See https://zqwang7.github.io/demos/ctPuLSE_demo/index.html.
@@ II Related Work
II-A Generalizability of Supervised Enhancement Models
@@
II-B Leveraging Close-Talk Mixtures for Speech Enhancement
Leveraging real-recorded close-talk mixtures as a weak supervision to train speech enhancement and speaker separation models directly on real-recorded far-field mixtures has attracted research interests recently. A representative algorithm in this direction is SuperM2M [3], which builds upon a weakly-supervised speaker separation algorithm named mixture to mixture (M2M) [24] and an unsupervised speaker separation algorithm named UNSSOR [25]. M2M trains enhancement models to produce a speech estimate and a noise estimate such that the two estimates can be linearly filtered to recover observed mixtures, and SuperM2M, building upon M2M, leverages supervised learning on simulated mixtures to improve M2M. Although SuperM2M has shown strong potential for neural speech enhancement and robust ASR, we often observe that SuperM2M cannot sufficiently suppress non-target signals, likely because its loss function is defined on reconstructed mixtures (i.e., the summation of linearly-filtered source estimates), rather than directly on source estimates. In comparison, the loss function in ctPuLSE, which we will show, is defined based on pseudo-labels derived from close-talk mixtures. As long as the pseudo-labels are high-quality, we can reasonably expect that the suppression of non-target signals would be more sufficient and aggressive.
@@ II-B Leveraging Close-Talk Mixtures for Speech Enhancement
II-C Pseudo-Label Based Speech Enhancement
@@
II-D Close-Talk Speech Enhancement
Our study leverages close-talk speech enhancement to derive high-quality pseudo-labels for far-field speech enhancement. There are existing studies [14, 15, 16] on close-talk speech enhancement, but for different purposes. For example, in [14, 15], the task is to enhance the target speech captured by a close-talk microphone with the help of an ear-mounted microphone; and in [16], the task is to enhance close-talk speech when users talk to dual-microphone mobile phones at a very short distance. Their purpose is not on training models on real data and improving the generalizability to real data.
@@ III Physical Model and Approach Overview
where the subscript $p\in\{1,\dots,P\}$ indexes the $P$ far-field microphones, subscript 00 indexes the close-talk microphone, $t\in\{0,\dots,T-1\}$ indexes $T$ frames, and $f\in\{0,\dots,F-1\}$ indexes $F$ frequency bins. In (1), $Y_{0}(t,f)$, $X_{0}(t,f)$, and $V_{0}(t,f)$ respectively denote the STFT coefficients of the close-talk mixture, close-talk speech, and non-speech signals (such as environmental noises) captured by the close-talk microphone at time $t$ and frequency $f$. Similarly, in (2), $Y_{p}(t,f)$, $X_{p}(t,f)$, and $V_{p}(t,f)$ are respectively the STFT coefficients of the far-field mixture, far-field speech, and non-speech signals captured by far-field microphone $p$ at time $t$ and frequency $f$. In the rest of this paper, when dropping the indices $p$, $t$ and $f$, we refer to the corresponding spectrograms. In this paper, $V$ is assumed to contain an unknown number of strong, non-stationary diffuse and directional noises sources.
@@ IV CTSEnet
where ||$\cdot$|| ⋅ | computes the magnitude or absolute value, $\mathcal{R}(\cdot)$ and $\mathcal{I}(\cdot)$ respectively extract the RI components, and the denominator in (3) balances the loss values across different training mixtures. Other configurations of CTSEnet are detailed later in Section VI-C and VI-D.
@@ V-A DNN Configurations
Next, we describe the loss functions for speech estimate $\hat{X}_{q}$ in Section V-B and for noise estimate $\hat{V}_{q}$ in V-D. Other DNN setups are detailed in Section VI-C and VI-D.
V-B Loss Functions Based on Estimated Close-Talk Speech
Assuming that the close-talk microphone and far-field microphones are approximately time-synchronized, we directly leverage $\hat{X}_{0}^{\text{CTSE}}$ as the pseudo-label for each far-field mixture $Y_{p}$. Since $\hat{X}_{0}^{\text{CTSE}}$ is time- and gain-aligned to close-talk speech $X_{0}$ rather than to far-field speech $X_{p}$ in far-field mixture $Y_{p}$, special care is needed to account for the time delay and gain differences between $\hat{X}_{0}^{\text{CTSE}}$ and $X_{p}$. In this context, we propose to first linearly filter the DNN estimate $\hat{X}_{p}$ at each frequency to align it to $\hat{X}_{0}^{\text{CTSE}}$ before loss computation:
@@ V-B Loss Functions Based on Estimated Close-Talk Speech
An alternative way for linear filtering is to estimate the filter in the time domain, following the ideas behind CI-SDR [35]. In detail, we compute a time-domain multi-tap Wiener filter to align $\hat{x}_{p}$ to $\hat{x}_{0}^{\text{CTSE}}$ (where $\hat{x}_{p}=\text{iSTFT}(\hat{X}_{p})$ is the re-synthesized time-domain signal of $\hat{X}_{p}$ obtained via inverse STFT (iSTFT), and similarly $\hat{x}_{0}^{\text{CTSE}}=\text{iSTFT}(\hat{X}_{0}^{\text{CTSE}})$) by solving the following problem:
@@
V-C Addressing Time-Synchronization Issues in Pre-Processiing
In the previous subseciton, when performing linear filtering for loss computation, we assume that far-field and close-talk microphones are reasonably time-synchronized. In practical data collection, the close-talk microphone and far-field microphone array are usually managed and processed by two different devices. Although the microphones on each device are typically synchronized, there could be synchronization errors between the microphones on different devices. Later in Section VI-B, we will introduce a technique based on the classic GCC-PHAT algorithm [36, 37] to roughly synchronize the close-talk microphone and far-field array. This technique is utilized as a pre-processing operation prior to training.
V-D Co-Learning on Simulated and Real Mixtures
In practical application scenarios, the amount of real data is typically scarce, as it is often effort-consuming to collect real-recorded close-talk and far-field mixture pairs. In this case, even if ctPuLSEnet can be trained on real data, the performance is often limited, simply due to the limited amount of training data. On the other hand, simulated data can be easily and massively generated through simulation. In this case, following the SuperM2M algorithm [3] which combines supervised learning on simulated data and weakly-supervised learning on real data, we propose to train ctPuLSEnet on both simulated and real data. See Fig. 3 for an illustration, where supervised learning on simulated data is shown in Fig. 3(b).
@@ V-D Co-Learning on Simulated and Real Mixtures
Alternatively, we can train ctPuLSEnet to not only predict target speech but also non-target signals via multi-task learning, as is described earlier in Section V-A. In this case, the loss on simulated data can be
@@
We can also add a weighting term $\alpha\in{\mathbb{R}}_{>0}$ between the loss on simulated data and the loss on real data to balance their importance. In detail, the overall loss is defined as
$\mathcal{L}_{q}=\left\{\begin{aligned} &\alpha\times\mathcal{L}_{q}^{\text{simu}},\text{for mini-batches of simulated data}\\ &\mathcal{L}_{q}^{\text{real}},\text{for mini-batches of real data}\end{aligned}\right.$ (14)
@@ VI Experimental Setup
Based on real-recorded mixtures, the main objective of our experiments is to demonstrate whether ctPuLSEnet can achieve better enhancement on real-recorded far-field mixtures than purely-supervised learning based approaches, which can only train enhancement models on simulated data. Bearing this objective in mind, we validate the proposed algorithms on the CHiME-4 dataset [40], which contains both real-recorded and simulated mixtures for both training and testing.222 Besides CHiME-4, we checked many other datasets for evaluation, but CHiME-4 is the only one with both real and simulated mixtures, paired close-talk and far-field mixtures, a setup for ASR evaluation, and a single target speaker in each recording. Closest to CHiME-4 are datasets such as CHiME-{5,6,7} [41, 10], AliMeeting [21] and AMI [42]. However, they are multi-speaker in nature and hence not very suitable for speech enhancement. Since, for real mixtures, the clean speech is not available for evaluation, we mainly check whether the enhanced speech can yield better ASR performance, considering that ASR scores can indicate the degrees of speech distortion and can reflect speech intelligibility. The ASR evaluation pipeline is shown in Fig. 4, where enhanced close-talk or far-field speech is directly fed to a strong pre-trained ASR system for recognition. The rest of this section describes the CHiME-4 dataset, synchronization of close-talk and far-field mixtures, setup for ASR evaluation, miscellaneous system configurations, baselines for comparison, and evaluation metrics.
VI-A CHiME-4 Dataset
@@
The mixtures are recorded in four representative daily environments (including buses, cafeteria, pedestrian areas and streets), where multiple strong, non-stationary directional and diffuse noises can naturally exist. In CHiME-4, the room reverberation is weak, and the major challenge is in how to deal with the multi-source non-stationary noise signals.
TABLE I: Number of Utterances in CHiME-4
Close-talk/
Type Far-field? #mics Training Set Validation Set Test Set
SIMU Far-field 6 $7,138$ ($\sim$15.115.115.115.1 h) $1,640$ ($\sim$2.92.92.92.9 h) $1,320$ ($\sim$2.32.32.32.3 h)
SIMU Close-talk - N/A N/A N/A
REAL Far-field 6 $1,600$ ($\sim$2.72.72.72.7 h) $1,640$ ($\sim$2.72.72.72.7 h) $1,320$ ($\sim$2.22.22.22.2 h)
REAL Close-talk 1 $1,600$ ($\sim$2.72.72.72.7 h) $1,640$ ($\sim$2.72.72.72.7 h) $1,320$ ($\sim$2.22.22.22.2 h)
The number of mixtures in CHiME-4 is summarized in Table I. Besides real mixtures, CHiME-4 also provides simulated far-field mixtures for training and testing. We emphasize that, for each real mixture in the training set, in total it has $7$ channels (i.e., $1$ close-talk plus $6$ far-field microphones), while, for each simulated training mixture, it is far-field, $6$-channel, and does not have the paired close-talk mixture.
@@
VI-B Synchronization of Close-talk and Far-field Mixtures
In CHiME-4, we observe that the far-field mixtures are reasonably synchronized with each other, while significant synchronization errors exist between close-talk and far-field mixtures. For some utterances, the time delay between the close-talk and far-field mixtures can be as large as $0.05$ seconds, which, if correct, means that the distance between the close-talk microphone and far-field array can be as large as $\sim$17171717 meters (i.e., $0.05\times 340$, assuming that the speed of sound in the air is $340$ meters per second). This clearly does not make sense in the CHiME-4 setup, as the speaker hand-holds the tablet while talking. The root cause of the synchronization errors, we think, is that the close-talk microphone and far-field array in CHiME-4 are placed on, and processed by, two different devices.
If the synchronization errors between the close-talk and far-field mixtures are not properly addressed before using them to train ctPuLSEnet, the loss functions in (5) would be less effective, since the hypothesized filter taps of $\mathbf{g}_{p}(f)$ in (7) or $h_{p}$ in (8), which are hyper-parameters shared by all the training utterances, may not be able to compensate the synchronization errors for every training utterance. To deal with this, we could use a very long linear filter to cover the maximum synchronization error of all the training mixtures. This is however problematic, as the longer the filter is, the more likely that the filter can filter any DNN estimate (even if the estimate is a random signal) to approximate the estimated close-talk speech.
To mitigate the synchronization errors, we design an algorithm, motivated by the classic GCC-PHAT algorithm [36, 37], to approximately align the close-talk mixture to far-field mixtures. We emphasize that this is a pre-processing step at the very beginning of our proposed system (i.e., before training).
In detail, for each of the close-talk and far-field mixtures, we consider its magnitude in each frequency as a one-dimensional sequence, and find an integer frame shift $\hat{d}\in{\mathbb{Z}}$ (shared by all the frequencies) that can, at every frequency, best align the magnitude sequence of the close-talk mixture to those of the far-field mixtures. Specifically, let $M_{p}(f)=|Y_{p}(\cdot,f)|\in{\mathbb{R}}^{T}$ denotes the magnitude sequence at frequency $f$ for microphone $p\in\{0,1,\dots,P\}$, and $R_{p}(f)=\text{FFT}\big{(}M_{p}(f)\big{)}\in{\mathbb{C}}^{T}$ denotes the FFT coefficients after applying a $T$-point faster Fourier transform (FFT) to the magnitude sequence. At each frequency $f$, we first compute the GCC-PHAT coefficients between the magnitude sequence of the close-talk microphone and that of a far-field microphone $p$ by
GCC-PHAT (${}_{p}(t,f,d)$

### v2

# ctPuLSE: Close-Talk, and Pseudo-Label Based Far-Field, Speech Enhancement

@@ Abstract
The current dominant approach for neural speech enhancement is via purely-supervised deep learning on simulated pairs of far-field noisy-reverberant speech (i.e., mixtures) and clean speech. The trained models, however, often exhibit limited generalizability to real-recorded mixtures. To deal with this, this paper investigates training enhancement models directly on real mixtures. However, a major difficulty challenging this approach is that, since the clean speech of real mixtures is unavailable, there lacks a good supervision for real mixtures. In this context, assuming that a training set consisting of real-recorded pairs of close-talk and far-field mixtures is available, we propose to address this difficulty via close-talk speech enhancement, where an enhancement model is first trained on simulated mixtures to enhance real-recorded close-talk mixtures and the estimated close-talk speech can then be utilized as a supervision (i.e., pseudo-label) for training far-field speech enhancement models directly on the paired real-recorded far-field mixtures. We name the proposed system ctPuLSE. Evaluation results on the popular CHiME-4 dataset show that ctPuLSE can derive high-quality pseudo-labels and yield far-field speech enhancement models with strong generalizability to real data.
preprint: Wang, JASA
@@ I Introduction
Dramatic progress has been made in speech enhancement [34], thanks to the rapid development of deep learning. The current dominant approach is based on supervised deep learning, where paired noisy-reverberant speech (i.e., mixtures) and clean speech are simulated and utilized to train deep neural networks (DNN) to predict the clean speech based on its paired mixtures in a purely-supervised, discriminative way [34, 53, 1]. The trained models, however, often exhibit limited generalizability to real-recorded mixtures [36, 24, 52, 30, 31, 10, 20, 9, 15, 51, 50], mainly because the simulated training data is typically, and in many cases inevitably, mismatched with real test data.
@@
Although ctPuLSE is simple and straightforward, it obtains strong robust ASR and speech enhancement performance on the public CHiME-4 dataset [2, 33, 3], the most popular benchmark to date in robust ASR and speech enhancement. The evaluation results suggest that ctPuLSE can effectively train enhancement models on real-recorded far-field mixtures, and can yield better generalizability to real data than purely-supervised models trained only on simulated data. A sound demo is provided at https://zqwang7.github.io/demos/ctPuLSE_demo/index.html.
@@ II Related Work
II.1 Generalizability of Supervised Enhancement Models
@@
II.2 Leveraging Close-Talk Mixtures for Speech Enhancement
Leveraging real-recorded close-talk mixtures as a weak supervision to train speech enhancement and speaker separation models directly on real-recorded far-field mixtures has attracted research interests recently. A representative algorithm in this direction is SuperM2M [36], which builds upon a weakly-supervised speaker separation algorithm named mixture to mixture (M2M) [35] and an unsupervised speaker separation algorithm named UNSSOR [43]. M2M trains enhancement models to produce a speech estimate and a noise estimate such that the two estimates can be linearly filtered to recover observed mixtures, and SuperM2M, building upon M2M, leverages supervised learning on simulated mixtures to improve M2M. Although SuperM2M has shown strong potential for neural speech enhancement and robust ASR, we often observe that SuperM2M cannot sufficiently suppress non-target signals, likely because its loss function is defined on reconstructed mixtures (i.e., the summation of linearly-filtered source estimates), rather than directly on source estimates. In comparison, the loss function in ctPuLSE, which we will show, is defined based on pseudo-labels derived from close-talk mixtures. As long as the pseudo-labels are high-quality, we can reasonably expect that the suppression of non-target signals by ctPuLSE would be more sufficient and aggressive.
@@ II.2 Leveraging Close-Talk Mixtures for Speech Enhancement
II.3 Pseudo-Label Based Speech Enhancement
@@
II.4 Relations to DAPS Dataset
A dataset named DAPS [23] is proposed for the task of transforming speech recorded on common consumer (and possibly band-limited) devices in real-world noisy-reverberant environments into professional production quality speech. A loud speaker is used to play back clean speech in real-world environments, and the speech is then recorded by consumer devices to create pairs of clean and degraded speech signals for model training. Although the DAPS paper proposes a task similar to the one in this study, it proposes a new dataset only rather than an actual algorithm. On the other hand, in our study the close-talk mixture is recorded by attaching a close-talk microphone to the target talker and hence would inevitably capture some interference signals, and the pseudo-label is then derived based on the close-talk mixture. Differently, DAPS uses played-back clean speech, which is used to derive pseudo-labels.
II.5 Close-Talk Speech Enhancement
Our study leverages close-talk speech enhancement to derive high-quality pseudo-labels for far-field speech enhancement. There are existing studies [16, 17, 29] on close-talk speech enhancement, but for different purposes. For example, in [16, 17], the task is to enhance the target speech captured by a close-talk microphone with the help of an ear-mounted microphone; and in [29], the task is to enhance close-talk speech when users talk to dual-microphone mobile phones at a very short distance. Their purpose is not on training speech enhancement models on real data and improving the generalizability to real data.
@@ III Physical Model and Approach Overview
where the subscript $p\in\{1,\dots,P\}$ indexes the $P$ far-field microphones, subscript 0 indexes the close-talk microphone, $t\in\{0,\dots,T-1\}$ indexes $T$ frames, and $f\in\{0,\dots,F-1\}$ indexes $F$ frequency bins. In (1), $Y_{0}(t,f)$, $X_{0}(t,f)$, and $V_{0}(t,f)$ respectively denote the STFT coefficients of the close-talk mixture, close-talk speech, and non-speech signals (such as environmental noises) captured by the close-talk microphone at time $t$ and frequency $f$. Similarly, in (2), $Y_{p}(t,f)$, $X_{p}(t,f)$, and $V_{p}(t,f)$ are respectively the STFT coefficients of the far-field mixture, far-field speech, and non-speech signals captured by far-field microphone $p$ at time $t$ and frequency $f$. In the rest of this paper, when dropping the indices $p$, $t$ and $f$, we refer to the corresponding spectrograms. In this paper, $V$ is assumed to contain an unknown number of strong, non-stationary diffuse and directional noises sources.
@@ IV CTSEnet
where $|\cdot|$ computes the magnitude or absolute value, $\mathcal{R}(\cdot)$ and $\mathcal{I}(\cdot)$ respectively extract the RI components, and the denominator in (3) balances the loss values across different training mixtures. Other configurations of CTSEnet are detailed later in Section VI.3 and VI.4.
@@ V.1 DNN Configurations
Next, we describe the loss functions for the speech estimate $\hat{X}_{q}$ in Section V.2 and for the noise estimate $\hat{V}_{q}$ in V.4. Other DNN setups are detailed in Section VI.3 and VI.4.
V.2 Loss Functions Based on Estimated Close-Talk Speech
Assuming that the close-talk microphone and far-field microphones are approximately time-synchronized, we directly leverage $\hat{X}_{0}^{\text{CTSE}}$ as the pseudo-label for each far-field mixture $Y_{p}$. Since $\hat{X}_{0}^{\text{CTSE}}$ is time- and gain-aligned to close-talk speech $X_{0}$ rather than to far-field speech $X_{p}$ in far-field mixture $Y_{p}$, special care is needed to account for the time delay and gain differences between $\hat{X}_{0}^{\text{CTSE}}$ and $X_{p}$.
In this context, we propose to first linearly filter the DNN estimate $\hat{X}_{p}$ at each frequency to align it to $\hat{X}_{0}^{\text{CTSE}}$ before loss computation:
@@ V.2 Loss Functions Based on Estimated Close-Talk Speech
An alternative way for linear filtering is to estimate the filter in the time domain, following the ideas behind CI-SDR [4]. In detail, we compute a time-domain multi-tap Wiener filter to align $\hat{x}_{p}$ to $\hat{x}_{0}^{\text{CTSE}}$, where $\hat{x}_{p}=\text{iSTFT}(\hat{X}_{p})$ is the re-synthesized time-domain signal of $\hat{X}_{p}$ obtained via inverse STFT (iSTFT), and similarly $\hat{x}_{0}^{\text{CTSE}}=\text{iSTFT}(\hat{X}_{0}^{\text{CTSE}})$. The filter is computed by solving the following problem:
@@
V.3 Addressing Time-Synchronization Issues in Pre-Processiing
In the previous subseciton, when performing linear filtering for loss computation, we assume that far-field and close-talk microphones are reasonably time-synchronized. In practical data collection, the close-talk microphone and far-field microphone array are usually managed and processed by two different devices. Although the microphones on each device are typically synchronized, there could be synchronization errors between the microphones on different devices. Later in Section VI.2, we will slightly modify the classic GCC-PHAT algorithm [11, 46] to roughly synchronize the close-talk microphone and far-field array. This technique is utilized as a pre-processing operation prior to training.
V.4 Co-Learning on Simulated and Real Mixtures
In practical application scenarios, the amount of real data is typically scarce, as it is often effort-consuming to collect real-recorded close-talk and far-field mixture pairs. In this case, even if ctPuLSEnet can be trained on real data, the performance is often limited, simply due to the limited amount of training data. On the other hand, simulated data can be easily and massively generated through simulation. In this case, following the SuperM2M algorithm [36], which combines supervised learning on simulated data with unsupervised or weakly-supervised learning on real data, we propose to train ctPuLSEnet on both simulated and real data. See Fig. 3 for an illustration, where supervised learning on simulated data is shown in Fig. 3(b).
@@ V.4 Co-Learning on Simulated and Real Mixtures
Alternatively, we can train ctPuLSEnet to not only predict target speech but also non-target signals via multi-task learning, as is described earlier in Section V.1. In this case, the loss on simulated data can be
@@
We can also add a weighting term $\alpha\in{\mathbb{R}}_{>0}$ between the loss on simulated data and the loss on real data to balance their importance. The overall loss is defined as
[…]
Table 1: Number of utterances in CHiME-4
Type Close-talk or far-field? #mics Training Set Validation Set Test Set
SIMU Far-field 6 $7,138$ ($\sim$$15.1$ h) $1,640$ ($\sim$$2.9$ h) $1,320$ ($\sim$$2.3$ h)
SIMU Close-talk - N/A N/A N/A
REAL Far-field 6 $1,600$ ($\sim$$2.7$ h) $1,640$ ($\sim$$2.7$ h) $1,320$ ($\sim$$2.2$ h)
REAL Close-talk 1 $1,600$ ($\sim$$2.7$ h) $1,640$ ($\sim$$2.7$ h) $1,320$ ($\sim$$2.2$ h)
@@ VI Experimental Setup
Based on real-recorded mixtures, the main objective of our experiments is to demonstrate whether ctPuLSEnet can achieve better enhancement on real-recorded far-field mixtures than purely-supervised learning based approaches, which can only train enhancement models on simulated data. Bearing this objective in mind, we validate the proposed algorithms on the CHiME-4 dataset [2, 33, 3], the most popular benchmark to date in robust ASR and speech enhancement. It contains both real-recorded and simulated mixtures for both training and testing. Since, for real mixtures, the clean speech is not available for evaluation, we mainly check whether the enhanced speech can yield better ASR performance, considering that ASR scores can reflect speech intelligibility and indicate the degrees of speech distortion. The ASR evaluation pipeline is shown in Fig. 4, where enhanced close-talk or far-field speech is directly fed to a strong pre-trained ASR system for recognition.
The rest of this section describes the CHiME-4 dataset, synchronization of close-talk and far-field mixtures, setup for ASR evaluation, miscellaneous system configurations, baselines for comparison, and evaluation metrics.
VI.1 CHiME-4 Dataset
@@
The mixtures are recorded in four representative daily environments (including buses, cafeteria, pedestrian areas and streets), where multiple strong, non-stationary directional and diffuse noises can naturally co-exist. In CHiME-4, the room reverberation is weak, and the major challenge is in how to deal with the multi-source non-stationary noise signals.
The number of mixtures in CHiME-4 is summarized in Table 1. Besides real mixtures, CHiME-4 also provides simulated far-field mixtures for training and testing. We emphasize that, for each real mixture in the training set, in total it has $7$ channels (i.e., $1$ close-talk plus $6$ far-field microphones), while, for each simulated training mixture, it is far-field, $6$-channel, and does not have the paired close-talk mixture.
@@
VI.2 Synchronization of Close-talk and Far-field Mixtures
In CHiME-4, we observe that the far-field mixtures are reasonably synchronized with each other, while significant synchronization errors exist between close-talk and far-field mixtures. For some utterances, the time delay between the close-talk and far-field mixtures can be as large as $0.05$ seconds, which, if correct, means that the distance between the close-talk microphone and far-field array can be as large as $\sim$$17$ meters (i.e., $0.05\times 340$, assuming that the speed of sound in the air is $340$ meters per second). This clearly does not make sense in the CHiME-4 setup, as the speaker hand-holds the tablet while talking. The root cause of the synchronization errors is that the close-talk microphone and far-field array in CHiME-4 are placed on, and processed by, two different devices.
If the synchronization errors between the close-talk and far-field mixtures are not properly addressed before using them to train ctPuLSEnet, the loss functions in (5) would be less effective, since the hypothesized filter taps of $\mathbf{g}_{p}(f)$ in (7) or $h_{p}$ in (8), which are hyper-parameters shared by all the training utterances, may not be able to compensate the synchronization errors for every training utterance. To deal with this, we could use a very long linear filter to cover the maximum synchronization error of all the training mixtures. This is however problematic, as the longer the filter is, the more likely that the filter can filter any DNN estimate (even if the estimate is a random signal) to accurately fit the estimated close-talk speech.
To mitigate the synchronization errors, we slightly modify the classic GCC-PHAT algorithm [11, 46] to approximately align the close-talk mixture to far-field mixtures. We emphasize that this is a pre-processing step at the very beginning of our proposed system (i.e., before training).
In detail, for each of the close-talk and far-field mixtures, we consider its magnitude in each frequency as a one-dimensional sequence, and find an integer frame shift $\hat{d}\in{\mathbb{Z}}$ (shared by all the frequencies) that can, at every frequency, best align the magnitude sequence of the close-talk mixture to those of the far-field mixtures. Specifically, let $M_{p}(f)=|Y_{p}(\cdot,f)|\in{\mathbb{R}}^{T}$ denote the magnitude sequence at frequency $f$ for microphone $p\in\{0,1,\dots,P\}$, and $R_{p}(f)=\text{FFT}\big(M_{p}(f)\big)\in{\mathbb{C}}^{T}$ denote the FFT coefficients after applying a $T$-point faster Fourier transform (FFT) to the magnitude sequence. At each frequency $f$, we first compute the GCC-PHAT coefficients between the magnitude sequence of the close-talk microphone and that of a far-field microphone $p$ by
GCC- $\text{PHAT}_{p}(t,f,d)$